uppose the level of output that creates full employment in the economy is 1 800. Using the multiplier, determine the level of investment spending that would create full employment in this economy


1
Expert's answer
2020-12-08T10:19:06-0500
"Solution"

The investment multiplier determines the economic impact of public or private investment. Mathematically, the investment multiplier is a function of Marginal Propensity to Consume(MPC) and the marginal Propensity to Save (MPS). An increase in average consumer spending ultimately leads to an increase in national output greater than the initial amount spent at a given MPC.Therefore;

Multiplier="\frac{1}{(1-MPC)}"

If the level of output is $1,800 and its MPC is 0.80, with an initial spending of 10M, then the new level of spending can be determined as below.

Step 1.

Determine the multiplier,

"\frac{1}{1-0.80}=5"

Then obtain the increase in spending, since the initial spending is 10M, and the multiplier is 5,then the increase in spending is "5\times10M=50M."

Since the initial output level is 1800M, then "1800+50=1850m" is the level that would create full employment in this economy.